Find the extreme values of f(x,y)=x+y^2 on the surface 2x^2+y^2=1.


1
Expert's answer
2021-06-08T17:47:12-0400

Ans:-



Solve equations "∇f= λ ∇g \ \ and \ \ \ g(x,y)=1" using Lagrange multipliers Constraint:





"g(x, y)=x^2+y^2=1"

Using Lagrange multipliers,





"f_x=\lambda g_x, f_y=\lambda g_y, g(x, y)=1"

which become





"2x=2x\lambda""1=2y\lambda""x^2+y^2=1"

If "x=0," then "y^2=1=>y=\pm1."

If "\lambda=1," then "y=\dfrac{1}{2}," and "x=\pm\dfrac{\sqrt{3}}{2}"

Therefore "f" has possible extreme values at the points "(0, -1), (0, 1),( -\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})," and "(\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})." Evaluating "f"at these four points, we find that 





"f(0, -1)=-1""f(0, 1)=1""f( -\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4}""f( \dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4}"

Therefore the maximum value of "f" on the circle "x^2+y^2=1" is





"f( \pm\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4},"

the minimum value of "f" on the circle "x^2+y^2=1" is





"f( 0,-1)=-1"